Simplfy: cos((3*)/4+x)+ sin((3*)/4-x)
Note: * = pie

Answer by RAY100(1637)

cos{ (3pi/4) +x} +sin { (3pi/4) -x}
.
The ref angle for (3pi/4) = pi/4
.
cos { (3pi/4 +x} = -cos (pi/4 -x),,,as long as x .
sin { (3pi/4)-x} = sin (pi/4 +x)
.
-cos (pi/4-x) + sin(pi/4 +x)
.
sin(pi/4 +x) - cos(pi/4-x)
.
Discussion: it is tempting to use the "sum and difference identities"
.
but this does not appear to "simplify"
.
Realizing that (pi/4) is 45 degrees, this is not a bad simplification